# Does Machine Unlearning Truly Remove Model Knowledge? A Framework for Auditing Unlearning in LLMs

### v1

@@ Abstract
In recent years, Large Language Models (LLMs) have achieved remarkable advancements, drawing significant attention from the research community. Their capabilities are largely attributed to large-scale architectures, which require extensive training on massive datasets. However, such datasets often contain sensitive or copyrighted content sourced from the public internet, raising concerns about data privacy and ownership. Regulatory frameworks, such as the General Data Protection Regulation (GDPR), grant individuals the right to request the removal of such sensitive information. This has motivated the development of machine unlearning algorithms that aim to remove specific knowledge from models without the need for costly retraining. Despite these advancements, evaluating the efficacy of unlearning algorithms remains a challenge due to the inherent complexity and generative nature of LLMs. In this work, we introduce a comprehensive auditing framework for unlearning evaluation, comprising three benchmark datasets, six unlearning algorithms, and five prompt-based auditing methods. By using various auditing algorithms, we evaluate the effectiveness and robustness of different unlearning strategies. To explore alternatives beyond prompt-based auditing, we propose a novel auditing technique based on intermediate activation perturbation. The complete framework will be open-sourced upon manuscript acceptance.
Does Machine Unlearning Truly Remove Model Knowledge? A Framework for Auditing Unlearning in LLMs
Haokun Chen1,2,3, Yueqi Zhang4, Yuan Bi4, Yao Zhang1,3, Tong Liu1,3, Jinhe Bi1,3, Jian Lan1,3, Jindong Gu5, Claudia Grosser2,4, Denis Krompass2, Nassir Navab4, Volker Tresp1,3 1 LMU Munich, Germany 2 Siemens AG, Munich, Germany 3 Munich Center for Machine Learning (MCML), Germany 4 Technical University of Munich, Germany 5 University of Oxford, UK
@@ 1 Introduction
Large language models (LLMs) have seen rapid advancements recently, resulting in improved performance and widespread adoption across numerous applications. These advancements are largely attributed to their large-scale architectures, which require training on datasets containing billions of tokens Kaplan et al. (2020). These datasets are typically constructed from large-scale corpora of publicly available internet text. However, such corpora often inadvertently include personally identifiable information (PII) or copyrighted material, which are considered sensitive and generally unsuitable for commercial use due to legal and ethical constraints. To comply with local regulations (e.g., GDPR li (2)) and internal organizational policies, it is often necessary to remove sensitive information from trained models.
Machine unlearning has emerged as a promising solution to this problem Cao and Yang (2015); Bourtoule et al. (2021a). This work is motivated by the legal framework proposed by the European Union, namely the GDPR li (2), which grants individuals the right to request the removal of their personal data in trained models. In particular, approximate unlearning seeks to remove specific knowledge from a model without the need for retraining from scratch Zhang et al. (2024a); Eldan and Russinovich (2023a); Izzo et al. (2021), while ensuring that the resulting model closely approximates a retrained counterpart within a bounded error. This approach is especially appealing in the context of LLMs, where full retraining is prohibitively expensive. Despite the development of numerous unlearning algorithms, few studies have systematically assessed their effectiveness and robustness Qi et al. (2024). Recent research has shown that many of these methods can be easily circumvented using simple paraphrasing attacks Shi et al. (2024).
@@
  • We propose an prompt-based auditing framework for evaluating unlearning in LLMs.
@@
2 Related Work
2.1 Machine Unlearning in LLMs
Machine unlearning has garnered significant attention in the context of LLMs. Various approaches for targeted knowledge removal have been proposed:  Eldan and Russinovich (2023a) removed Harry Potter-related knowledge by finetuning LLMs on corpora with replaced keywords.  Zhang et al. (2024b) proposed to steer model preferences in the negative direction to reduce memorization.  Wang et al. (2024) used reversed knowledge distillation to eliminate personal information.  Feng et al. (2024) introduced a reweighted gradient ascent method for unlearning, and  Pawelczyk et al. (2023) utilized in-context unlearning examples.  Liu et al. (2024a); Bhaila et al. (2024) adapted input embeddings associated with the unlearning target, while Li et al. (2024); Tamirisa et al. (2024); Huu-Tien et al. (2024); Ashuach et al. (2024) proposed interventions in the model activation space. From a safety perspective, several studies have targeted the unlearning of harmful or undesirable model behaviors Zhang et al. (2024c); Yao et al. (2023); Liu et al. (2024c), while others have addressed backdoor unlearning Liang et al. (2024); Liu et al. (2024b) to enhance model integrity.
2.2 Evaluation of LLM Unlearning
While advances have been made in developing machine unlearning algorithms for LLMs, rigorous methodologies for auditing the efficacy of unlearning remain understudied. Recent work has demonstrated that unlearned knowledge can be recovered using techniques such as adversarial soft token attacks Schwinn et al. (2024) or few-shot in-context prompting Doshi and Stickland (2024). When model parameters are accessible, recovery can also be achieved through model quantization Zhang et al. (2024d) or retraining on partially unlearned datasets Łucki et al. (2024); Hu et al. (2024).  Lynch et al. (2024) compared eight unlearning evaluation methods and found persistent similarity in the latent representations of unlearned models.  Shi et al. (2024) conducted a comprehensive analysis across six evaluation dimensions using news and book datasets, highlighting the fragility of unlearning to paraphrasing and unrelated data.  Qi et al. (2024) showed that fine-tuning on unrelated data could inadvertently restore previously removed knowledge. Several benchmarks have been developed to evaluate the existing unlearning algorithms.  Maini et al. (2024) proposed a benchmark centered on fictitious author profiles, while Jin et al. (2024) introduced the Real-World Knowledge Unlearning (RWKU) benchmark, which targets knowledge removal of 200 well-known individuals. Complementary to this, the WDMP benchmark Li et al. (2024) focuses on the unlearning of sensitive knowledge related to biosecurity and cybersecurity threats.
Figure 1: The proposed activation perturbation-based algorithm (ActPert) for auditing unlearning in LLMs.
@@ 3 Preliminaries
Large Language Models (LLMs) operate within an auto-regressive framework, generating text by predicting the next token in a sequence based on preceding tokens. Given an input sequence of tokens $x_{1:t}$, the model computes the conditional probability distribution $p(x_{t+1}|x_{1:t})$ over the vocabulary at each time step. The overall likelihood of the sequence is given by:
$\log p(x_{t+1}|x_{1:t})=\sum_{t=1}^{T}\log p(x_{t}|x_{1:t-1})$ (1)
During inference, tokens are generated iteratively: at each step, the model samples the next token $x_{t+1}$ from the predicted distribution (using methods such as greedy decoding or nucleus sampling Holtzman et al. (2019)), then appends it to the existing context $x_{1:t}$ for the subsequent prediction.
Machine unlearning refers to the process of removing the influence of specific data from a trained model. Consider a machine learning model $f$ trained on a dataset $D_{train}$. When a data owner requests the removal of a subset $D_{u}\in D_{train}$, the goal of machine unlearning is to produce a modified model $f_{u}$ that behaves as if it had never been trained on $D_{u}$. Unlearning techniques generally fall into two categories: exact unlearning, which seeks to fully eliminate the impact of the forgotten data, and approximate unlearning, which aims for partial or probabilistic removal. We introduce them in the following.
Exact unlearning guarantees that the output distribution of the unlearned model $f_{u}$ is statistically indistinguishable from that of a reference model $f_{ref}$, which is trained from scratch on the retained dataset $D_{r}=D_{train}-D_{u}$. This strong guarantee of data removal is formalized as:
$p(f_{u}(x)=y)$ $=p(f_{ref}(x)=y)\quad$ (2)
$s.t.\quad\forall(x,y)\in D_{train}.$
While retraining from scratch is the most direct method to achieve exact unlearning, it is often computationally infeasible for large-scale models such as LLMs. One solution is to partition $D_{train}$ into overlapping subsets and train an ensemble of models Bourtoule et al. (2021b). Upon receiving an unlearning request, only the models that contain the targeted records need to be retrained.
@@
In the context of LLMs, approximate unlearning is typically realized through information overwriting Eldan and Russinovich (2023a); Wang et al. (2024), behavioral steering Feng et al. (2024), or model editing—via weight or activation modifications Liu et al. (2024a); Bhaila et al. (2024); Li et al. (2024); Tamirisa et al. (2024); Huu-Tien et al. (2024); Ashuach et al. (2024); Meng et al. (2022a, b). These methods aim to diminish or redirect the model’s reliance on the forgotten data without necessitating a full retraining cycle.
@@ 4 Proposed Method
Before introducing the proposed framework for unlearning auditing, we introduce Activation Perturbation-based Auditing (ActPert), a method for probing residual knowledge in unlearned language models. A schematic overview of the proposed approach is shown in Figure 1. Our method builds on recent advances in activation engineering for LLMs Arditi et al. (2024), which compute “refusal directions” by contrasting activations between harmful and harmless prompts to reduce a model’s tendency to refuse answering.
Analogously, we treat unlearning targets as sensitive queries and seek to perturb their input representations such that they become effectively insensitive, thereby increasing the chances of eliciting meaningful responses. Concretely, we inject random noise into the token embeddings corresponding to the unlearning target (e.g., the phrase Harry Potter in the prompt "Who are Harry Potter’s two best friends?"). This noise injection prevents the model from directly attending to the sensitive content during inference, resulting in a set of $n_{p}$ perturbed versions of the original query in the embedding space.
@@ 5 Experiments and Analyses
In this section, we provide details about the proposed unlearning auditing framework, which encompasses commonly used unlearning algorithms as well as established benchmarks. We begin by outlining representative benchmarks, unlearning methods, and auditing methods used in our evaluation. Following this, we report validation results from multiple auditing algorithms, including our proposed ActPert, and provide a comparative analysis to assess their effectiveness in detecting residual knowledge.
@@ 5.1 Unlearning Benchmarks
  • WHP Eldan and Russinovich (2023b). We audit a pretrained language model that was finetuned to unlearn knowledge related to Harry Potter. As the authors only released the unlearned model checkpoint without the original unlearning dataset $D_{u}$, we use GPT-4o to generate Q&A pairs for auditing. Specifically, we prompt GPT-4o to produce factual, short-form questions about Harry Potter. These generated questions are then submitted to both the pretrained and the unlearned models. Questions for which both models produce either correct or incorrect answers are filtered out. Further details about the filtering process are provided in the appendix. In total, we select 35 Q&A pairs for auditing.
  • TOFU Maini et al. (2024). TOFU is an unlearning benchmark constructed from autobiographies of 200 fictitious authors generated using GPT-4o, ensuring no prior presence in LLM pretraining. The pretrained language model is first finetuned on the full dataset, then unlearning is applied to either 1% (2 authors) or 5% (10 authors) of the training data. We follow the original paper’s hyperparameter settings for both fine-tuning and unlearning. While the original TOFU dataset consists of long-form Q&A pairs, we adopt GPT-4o to generate short-form questions and maintain consistency with other benchmarks. We filter out questions that the finetuned model answers incorrectly. This process yields 16 Q&A pairs for auditing the model unlearned with 1% of TOFU and 80 Q&A pairs for the model unlearned with 5%.
  • RWKU Jin et al. (2024). RWKU is a benchmark focused on real-world knowledge unlearning. It constructs Q&A pairs about well-known public figures as unlearning targets. This approach avoids the need for retraining, as such knowledge is prevalent in open-source LLMs. The unlearning dataset $D_{u}$ consists of biographical texts about these individuals, and RWKU includes Q&A pairs for evaluating different levels of unlearning effectiveness. To examine the impact of unlearning dataset size, we select 10, 20, and 30 individuals as unlearning targets and evaluate all unlearned models using the same 10 individuals.
@@ 5.3 Unlearning Algorithms
  • Gradient Ascent (GA) minimizes the probability that the target model $f_{u}$ makes correct predictions on the unlearning set $D_{u}$.
    $\mathcal{L}_{GA}=\frac{1}{|D_{u}|}\sum_{x\in D_{u}}\ell(x,f_{u}).$ (4)
  • Gradient Difference (GD) is a variant of GA that incorporates an additional loss term to preserve performance on the retain set $D_{r}$.
    $\mathcal{L}_{GD}=-\frac{1}{|D_{r}|}\sum_{x\in D_{r}}\ell(x,f_{u})+\mathcal{L}_{GA}(D_{u},f_{u}).$ (5)
  • Knolwedge Distillation (KD) extends GA by minimizing the KL divergence between the output token probabilities of the fine-tuned model ($f_{ft}$) and the unlearned model on the retained dataset $D_{r}$.
    $\begin{split}\mathcal{L}_{KD}=&-\frac{1}{|D_{r}|}\sum_{x\in D_{r}}\mathcal{L}_{KL}(f_{ft}(x)||f_{u}(x))\\ &+\mathcal{L}_{GA}(D_{u},f_{u}).\end{split}$ (6)
  • Rejection Tuning (RT) Maini et al. (2024) aligns the model to refuse when queried about target knowledge. This is achieved by constructing $D_{u}^{\text{idk}}$, where the responses to questions in the unlearning set $D_{u}$ are replaced with I don’t know or similar refusal-style responses.
    $\begin{split}\mathcal{L}_{RT}=&-\frac{1}{|D_{r}|}\sum_{x\in D_{r}}\ell(x,f_{u})\\ &-\frac{1}{|D_{u}^{idk}|}\sum_{x\in D_{u}^{idk}}\ell(x,f_{u}).\end{split}$ (7)
  • Direct Preference Optimization (DPO) (Rafailov et al., 2023) aligns the model to avoid providing accurate target knowledge. Specifically, counterfactual responses are used as positive examples $y_{w}$, consisting of model-generated fabrications about the unlearning target, while negative responses $y_{l}$ are drawn based on the ground-truth knowledge:
    $\begin{split}\mathcal{L}_{DPO}=&-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{u}}[\log\sigma\\ &(\beta\log\frac{f_{u}(y_{w}\mid x)}{f_{ft}(y_{w}\mid x)}-\beta\log\frac{f_{u}(y_{l}\mid x)}{f_{ft}(y_{l}\mid x)})]\end{split}$ (8)
  • Negative Preference Optimization (NPO) (Zhang et al., 2024b) is a variant of DPO that retains only the negative examples, i.e., the ground-truth knowledge to be unlearned, without any positive samples.
    $\mathcal{L}_{NPO}=-\mathbb{E}_{(x,y_{l})\sim\mathcal{D}_{u}}[\log\sigma(-\beta\log\frac{f_{u}(y_{l}\mid x)}{f_{ft}(y_{l}\mid x)})]$ (9)
@@ 5.4 Auditing Algorithms
  • AOA (Liu et al., 2023) adds a prefix that prompts the LLM to role-play as an Absolutely Obedient Agent, ensuring it strictly follows user instructions without deviation.
@@
  • GCG (Zou et al., 2023) optimizes an adversarial suffix to compel the model to produce affirmative responses, such as "Sure, the answer is…", instead of refusals. GCG was originally designed to jailbreak LLMs and generate responses to harmful queries, while we adapt it in our study to audit unlearned models.
  • SoftGCG (Schwinn et al., 2024) is a variant of GCG that optimizes the adversarial suffix in the token embedding space, enabling gradient-based optimization and improving attack success rates.
To enhance the thoroughness of the experiment, we combine various instruction-based auditing algorithms and present the corresponding results.
@@ 5.5.1 Main Results Analysis
The auditing performance of various algorithms is presented in Table 1. We utilize the ROUGE score to evaluate auditing performance. Specifically, we report the length of the longest common subsequence (LCS) between the reference answer and the model response, normalized by the length of the reference response. We include the validation performance of the pretrained (for RWKU and WHP) or finetuned (for TOFU) models (shown in brackets), the unlearned model (base), and the results of applying different auditing methods to the unlearned model.
Overall, a notable performance gap is observed between the baseline (no auditing) and the application of different auditing methods, indicating that residual knowledge persists in the model after unlearning. Among the approaches, prefix-based auditing methods—such as AOA—generally outperform optimization-based methods like SoftGCG. This suggests that modifying the query prompt alone can already prompt correct responses, thereby revealing latent retained knowledge. Although our proposed method ActPert does not consistently surpass all baselines across every scenario, it demonstrates competitive or superior results in several settings, particularly with GA and NPO. We prodivide detailed analysis for different benchmarks in the following.
On the RWKU benchmark, we observe that DPO and GA outperform NPO in terms of unlearning performance, both with and without auditing. Notably, RT achieves the best results without auditing, but its performance deteriorates significantly when an auditing method is applied. This indicates that rejection-based approaches, may not be as robust as removal-based methods for effective unlearning. Additionally, as the size of the unlearning dataset increases, the performance of GA and DPO tends to improve, while methods like NPO remain relatively stable. From an architectural perspective, the Phi-3 model exhibits greater resilience to auditing attacks compared to the LLaMA model.
For the TOFU dataset, preference-based unlearning algorithms generally perform better than other unlearning methods. Moreover, larger unlearning datasets tend to enhance unlearning effectiveness. Notably, unlearning efficacy on TOFU is significantly higher than on RWKU, and the application of auditing algorithms has minimal impact on the results. We hypothesize that this is due to TOFU being constructed using fictitious author biographies, which allows direct access to ground-truth queries during machine unlearning. However, this advantage is not typically available in most real-world applications.
On the WHP dataset, AOA stands out as the most effective auditing method. Therefore, we assume that the keyword-replacement-based unlearning algorithms used in WHP benchmark appears less effective compared with other methods.

### v2

# Lock-LLM: Prevent Unauthorized Knowledge Use from LLMs

@@ Abstract
In recent years, Large Language Models (LLMs) have achieved remarkable advancements, drawing significant attention from the research community. Their capabilities are largely attributed to large-scale architectures, which require extensive training on massive datasets. However, such datasets often contain sensitive or copyrighted content sourced from the public internet, raising concerns about data privacy and ownership. Regulatory frameworks, such as the General Data Protection Regulation (GDPR), grant individuals the right to request the removal of such sensitive information. This has motivated the development of machine unlearning algorithms that aim to remove specific knowledge from models without the need for costly retraining. Despite these advancements, evaluating the efficacy of unlearning algorithms remains a challenge due to the inherent complexity and generative nature of LLMs. In this work, we introduce a comprehensive auditing framework for unlearning evaluation, comprising 3 benchmark datasets, 6 unlearning algorithms, and 5 prompt-based auditing methods. By using various auditing algorithms, we evaluate the effectiveness and robustness of different unlearning strategies. To explore alternatives beyond prompt-based auditing, we propose a novel auditing technique based on intermediate activation perturbation. This approach offers a new perspective and serves as a potential direction for the future design of auditing algorithms. The complete framework and the proposed algorithm will be open-sourced upon manuscript acceptance.
@@ 1 Introduction
Large language models (LLMs) have seen rapid advancements recently, resulting in improved performance and widespread adoption across numerous applications. These advancements are largely attributed to their large-scale architectures, which require training on datasets containing billions of tokens [15]. These datasets are typically constructed from large-scale corpora of publicly available internet text. However, such corpora often inadvertently include personally identifiable information (PII) or copyrighted material, which are considered sensitive and generally unsuitable for commercial use due to legal and ethical constraints. To comply with local regulations (e.g., GDPR [16]) and internal policies, it is often necessary to remove sensitive information from trained models.
Machine unlearning has emerged as a promising solution to this problem [5, 4]. This work is motivated by the legal framework proposed by the European Union, namely the GDPR [16], which grants individuals the right to request the removal of their personal data in trained models. In particular, approximate unlearning seeks to remove specific knowledge from a model without the need for retraining from scratch [37, 7, 13], while ensuring that the resulting model closely approximates a retrained counterpart within a bounded error. This approach is especially appealing in the context of LLMs, where full retraining is prohibitively expensive. Despite the development of numerous unlearning algorithms, few studies have systematically assessed their effectiveness and robustness [30]. Recent research has shown that many of these methods can be easily circumvented using simple paraphrasing attacks [33].
@@
  • We propose a prompt-based auditing framework for evaluating unlearning in LLMs.
@@ 2 Preliminaries
Machine unlearning refers to the process of removing the influence of specific data from a trained model. Consider a machine learning model $f$ trained on a dataset $D_{train}$. When a data owner requests the removal of a subset $D_{u}\in D_{train}$, the goal of machine unlearning is to produce a modified model $f_{u}$ that behaves as if it had never been trained on $D_{u}$. Unlearning techniques generally fall into two categories: exact unlearning, which seeks to fully eliminate the impact of the forgotten data, and approximate unlearning, which aims for partial or probabilistic removal.
While retraining from scratch is the most direct method to achieve exact unlearning, it is often computationally infeasible for large-scale models such as LLMs. Therefore, we focus on approximate unlearning in this work.
@@
In the context of LLMs, approximate unlearning is typically realized through information overwriting [7, 35], behavioral steering [9], or model editing—via weight or activation modifications [20, 3, 17, 34, 12, 2, 27, 28]. These methods aim to diminish or redirect the model’s reliance on the forgotten data without necessitating a full retraining cycle.
@@ 3 Proposed Method
Before introducing the proposed framework for unlearning auditing, we introduce Activation Perturbation-based Auditing (ActPert), a method for probing residual knowledge in unlearned language models. A schematic overview of the proposed approach is shown in Figure 1. Our method builds on recent advances in activation engineering for LLMs [1], which compute “refusal directions” by contrasting activations between harmful and harmless prompts to reduce a model’s tendency to refuse answering.
Analogously, we treat unlearning targets as sensitive queries and seek to perturb their input representations such that they become effectively insensitive, thereby increasing the chances of eliciting meaningful responses. Concretely, we inject random noise into the token embeddings corresponding to the unlearning target (e.g., the phrase Harry Potter in the prompt "Who are Harry Potter’s two best friends?"). This noise injection prevents the model from directly attending to the sensitive content during inference, resulting in a set of $n_{p}$ perturbed embedding of the original query.
@@ 4 Experiments and Analyses
In this section, we provide details about the proposed unlearning auditing framework, which encompasses commonly used unlearning algorithms as well as established benchmarks. We begin by outlining benchmarks, unlearning methods, and auditing methods in our evaluation. Following this, we report validation results from multiple auditing algorithms, including our proposed ActPert, and provide a comparative analysis to assess their effectiveness in detecting residual knowledge.
@@ 4.1 Unlearning Benchmarks
  • WHP [8]. We audit the model finetuned to unlearn Harry Potter knowledge. Since the unlearning dataset $D_{u}$ is unavailable, we generate 35 short factual Q&A pairs with GPT-4o, of which pairs both pretrained and unlearned models answer identically are filtered out. Further details about the filtering process are provided in the Appendix.
  • TOFU [26]. TOFU uses autobiographies of 200 fictitious authors created with GPT-4o. Following the original setup, the model is finetuned on the full dataset and then unlearned on 1% (2 authors) or 5% (10 authors). We generate short questions with GPT-4o and discard pairs the finetuned model fails to answer. This yields 16 Q&As (1%) and 80 Q&As (5%).
  • RWKU [14]. RWKU targets real-world knowledge by unlearning facts about public figures. $D_{u}$ consists of biographical texts, with Q&A pairs used to measure unlearning effectiveness. To study dataset size effects, we unlearn 10, 20, and 30 individuals and evaluate all models on the same 10-person subset.
@@ 4.3 Unlearning Algorithms
  • Gradient Ascent (GA) [26] minimizes the probability that the target model $f_{u}$ makes correct predictions on the unlearning set $D_{u}$.
  • Gradient Difference (GD) [19] is a variant of GA that incorporates an additional loss term to preserve performance on the retain set $D_{r}$.
  • Knowledge Distillation (KD) [10] extends GA by minimizing the KL divergence between the output token probabilities of the fine-tuned model ($f_{ft}$) and the unlearned model on the retained dataset $D_{r}$.
  • Rejection Tuning (RT) [26] aligns the model to refuse when queried about target knowledge. This is achieved by constructing $D_{u}^{\text{idk}}$, where the responses to questions in the unlearning set $D_{u}$ are replaced with I don’t know or similar refusal-style responses.
  • Direct Preference Optimization (DPO) [31] aligns the model to suppress accurate target knowledge by using fabricated counterfactual responses as positives $y_{w}$ and ground-truth answers as negatives $y_{l}$.
  • Negative Preference Optimization (NPO) [38] is a DPO variant that retains only the ground-truth knowledge to be unlearned as negatives.
@@ 4.4 Auditing Algorithms
  • AOA [22] adds a prefix that prompts the LLM to role-play as an Absolutely Obedient Agent, ensuring it strictly follows user instructions without deviation.
@@
  • GCG [41] optimizes an adversarial suffix to compel the model to produce affirmative responses, such as "Sure, the answer is…", instead of refusals. GCG was originally designed to jailbreak LLMs and generate responses to harmful queries, while we adapt it in our study to audit unlearned models.
  • SoftGCG [32] is a variant of GCG that optimizes the adversarial suffix in the token embedding space, enabling gradient-based optimization and improving attack success rates.
@@ 4.5.1 Main Results Analysis
The auditing results are reported in Table 1, measured with ROUGE-L (longest common subsequence normalized by reference length). We include validation scores for the pretrained (RWKU, WHP) or finetuned (TOFU) models [in brackets], the unlearned baseline (base), and the outcomes of different auditing methods applied to the unlearned models.
Across benchmarks, a consistent gap emerges between the unlearned baseline and the audited models, indicating that residual knowledge remains accessible after unlearning. Prefix-based auditors such as AOA generally outperform optimization-based methods like SoftGCG, showing that prompt manipulation alone can reveal suppressed knowledge. Our proposed method ActPert delivers competitive performance and in several cases surpasses baselines, particularly when paired with GA and NPO.
On RWKU, DPO and GA achieve stronger unlearning than NPO, both with and without auditing. RT attains the best baseline score but degrades sharply once audited, suggesting that rejection-based strategies are less robust than removal-based methods. Larger unlearning datasets further improve GA and DPO, whereas NPO remains relatively unaffected. Among architectures, Phi-3 is more resistant to auditing than LLaMA, underscoring model-dependent resilience.
For TOFU, preference-based algorithms consistently outperform others, and larger unlearning sets enhance performance. Unlearning on TOFU is markedly more effective than on RWKU, and auditing has limited influence. This likely stems from TOFU’s synthetic construction, where fictitious biographies provide direct ground-truth queries—an advantage rarely present in real-world applications.
On WHP, AOA proves the most effective auditor, exposing the limitations of keyword-replacement unlearning approaches. This suggests that methods tailored to specific patterns may not generalize well to diverse or adversarial queries.